OpenStudy (anonymous):

For the below pair of group show that they are not isomorphic or construct explicitly the isomorphisms between them: (R,+) and the set of 2x2 matrices (1 x 0 1) where x is a real number, under matrix multiplication

OpenStudy (jamesj):

What's your intuitive sense on this?

OpenStudy (anonymous):

They are both infinite sets so I guess they are not isomorphic

OpenStudy (jamesj):

Two infinite groups can be isomorphic, so that's not a good argument.

OpenStudy (jamesj):

Let me ask you, what is the product of two arbitrary members of the matrix group?

OpenStudy (jamesj):

I.e., what is the product of 1 x 0 1 and 1 y 0 1

OpenStudy (anonymous):

1 x+y 0 1

OpenStudy (anonymous):

This group theory is so confusing for me :(

OpenStudy (jamesj):

Yes, that matrix multiplication is correct. And that suggests there might be a very natural isomorphism between (R,+) and this matrix group, call it (M, .)

OpenStudy (anonymous):

It looks just like addition but one is just a number the other is a matrix. So I can see something common but dont really get it

OpenStudy (anonymous):

Am I right that we are searching for a function that maps from one group to the other? But how is that possible between 2 different things?

OpenStudy (jamesj):

Define a function \[ f : (\mathbb{R},+) \rightarrow (M, .) \] by \[f(x) = \left[\begin{matrix}1 & x \\ 0 & 1\end{matrix}\right]\]

OpenStudy (jamesj):

Now you just need to prove that is an isomorphism.

OpenStudy (anonymous):

Things like that it is closed?

OpenStudy (jamesj):

Actually, first you need to show that it is in fact a group homomorphism: i.e., f(x + y) = f(x).f(y), f(0) = 0

OpenStudy (jamesj):

and then you need to show it is 1:1 and onto

OpenStudy (jamesj):

"Things like it is closed" is proved that the group operation doesn't take two elements of a set to something outside the set, so not a property of group homomorphisms.

OpenStudy (anonymous):

Yes, Im just reading my notes, that was something else. I found this: have to have the same size (that is cool) Have to have the same order. (what is the order of these?)

OpenStudy (jamesj):

Both have infinite order, because for every non-identity member x in (R,+) and m in M, there is no integer n such that nx = 0 or nm = 0.

OpenStudy (jamesj):

Those are necessary conditions, but not sufficient.

OpenStudy (anonymous):

I found it! Has to be a bijection and alfa(x)alfa(y)=alfa(xy)

OpenStudy (anonymous):

OK I will give it a go to prove the bijection but Im absolutely not sure if I can do it

OpenStudy (jamesj):

right, that's the function I wrote down. note that the group operation in (R,+) is addition, +; but in (M, .), the group operation is matrix multiplication. So you need to show that f(x+y) = f(x)f(y), where f(x)f(y) is the product of the two matrices f(x), f(y).

OpenStudy (anonymous):

What I did: f(x)=f(y) that is (1 x = 1 y 0 1 0 1 Implies that x=y so its 1:1 f(x+y)=(1 x+y = (1 x ( 1 y = f(x) f(y) 0 1 ) 0 1) 0 1)

OpenStudy (anonymous):

What do I do with the onto part?

OpenStudy (jamesj):

You need to show that for an arbitrary member m of M, there is a g in R such that f(g) = m.

OpenStudy (anonymous):

but that is trivial

OpenStudy (jamesj):

Indeed. :-)

OpenStudy (anonymous):

Thanks a lot, I have a few more of these. I will struggle with them for a while and if Im stuck I post a new question

OpenStudy (jamesj):

ok